OpenStudy (anonymous):

algebra problem: 5 - 1/2(4x + 6) = -3x -8

OpenStudy (anonymous):

\[5 - _{2}^{1}(4x + 6) = -3x - 8\] how do I solve this? I know the answer is -10 but I can't get it using distributive property.

OpenStudy (anonymous):

you want to find x right? :)

OpenStudy (anonymous):

Let's take it one step at a time ^_^: 5-2x-3 = -3x -8 now let's get all the x's on one side and the values on the other :) : -2x+3x = -8 -5 + 3 x= -10 Hope it's clearer now ^_^